# A causal inference framework for spatial confounding

### v11

@@ 4 Existing approaches to spatial confounding
Splines can be used to model $g$. Alternatively, a linear mixed model (LMM)
@@
is used with the random effects $w(s_{i})$ typically modeled with Gaussian process (GP), i.e., $w$ is a random function of space. The spatial effect is assumed to fully capture any spatial dependence, leaving the errors $\epsilon$ to be iid.
In what follows in this section, we assume (3) as the data generation model while the analysis model can be (3) or (4) or other choices. In most applications, there is nothing inherently random about the true functions or surfaces modeled using random effects. For example, using a GP to analyze topographical elevation data does not commit the researcher to the view that elevation is a random quantity. Also, assuming data to be generated from a random effect model may lead to spurious claims of unbiasedness of certain estimators as we illustrate below.
@@
Ordinary least squares: The naive, unadjusted estimate of the effect of $X$ on $Y$ is given by the ordinary least squares (OLS) estimator, regressing $Y$ on $X$. This is the baseline to which spatial confounding proposals are often compared. It is unbiased assuming the random effect model (4) as the true data generation process, but see Section S4 of the supplement for a toy illustration of how claims of unbiasedness of the OLS estimator can be misleading. When data are generated from a process like (3) with a spatial effect, it is well-known that this procedure does not adjust for spatial confounding–under any of the operative definitions thereof–and is biased in most spatial settings (see Dupont et al., 2020, for a discussion in the context of spatial confounding).
@@
Gaussian Process regression: The LMM analysis model (4) with a GP prior for the random function $w$ calculates probabilities under the assumption that the random effect is independent of the exposure. At first glance, this would seem to rule out the possibility of addressing causal confounding. However, LMMs (4) with GP prior can consistently estimate $\beta$ when the data are generated from a PLM (Bickel and Kleijn, 2012; Yang et al., 2015). Therefore both spline models and GP regression are valid analysis models for estimating the effect of exposure when the data are generated from (3) and (5).
@@
Restricted spatial regression: Restricted spatial regression (RSR) (Hodges and Reich, 2010), considers the same LMM as (4) but restricts $w$ to be orthogonal to the exposure $X$. Thus, including the spatial effects does not change the fixed effect estimate and $\bm{\hat{\beta}}_{RSR}=\bm{\hat{\beta}}_{OLS}$. The fact that the estimates from the mixed-effects model, like the GP regression model or the GLS estimator obtained from marginalizing out the LMM, are different from the OLS estimate was seen as a problem by Hodges and Reich (2010). This overlooks the fact that OLS is inconsistent for the parameter of a PLM, as discussed above. The viewpoint of Hodges and Reich (2010) can perhaps be seen as stemming from assuming the LMM (4) as the data-generating process, in which case the OLS is consistent. The GLS estimate, accounting for spatial structure via the error term, will differ from the OLS estimate in finite samples, but both are unbiased estimators in this scenario. Even assuming (4) as the true data-generating process, RSR has been criticized by Khan and Calder (2020) and Zimmerman and Hoef (2021) as producing anti-conservative intervals. However, we emphasize that with respect to the issue of causal confounding, RSR does not even attempt to supply an unbiased effect estimate; rather, it assumes there is no confounding bias.
@@
Spatially varying coefficient model (SVC): Gelfand et al. (2003) maintains the assumption of a linear outcome regression as in (3) or (4) but allows the slope of the regression to vary by location, i.e., $\beta=\beta(S_{i})$. This SVC model thus relaxes the assumption of effect homogeneity and can account for confounding under spatially heterogeneous effects.
@@
Limitations of linear models: Under the assumption that data are generated from a partially linear model (3) (or an SVC) and the exposures are generated from (5), the spline- and GP-based methods above can be used to estimate causal effects under spatial confounding. To see why, we note that the assumption of (3) and (5) as the true data generation process implies the causal identification assumptions proposed in Section 3, and thus that controlling for a flexible function of space suffices to control for unmeasured spatial confounding. This is stated formally in the proposition below. The proof is provided in the Supplementary Section S1.
Proposition 4.1.
@@
Recent work in spatial statistics extended the GP regression to allow for non-linear covariate effects fitted flexibly using random forests or boosting, thereby creating GLS versions of these machine learning procedures (Saha et al., 2021; Sigrist, 2020). These extensions still use additive models, implying that the effect of exposure is homogeneous in space. In Section 5, we propose a class of methods that simultaneously relax both the linearity and homogeneity assumptions.
4.1 Recent causal advances
Papadogeorgou et al. (2018) extend propensity score-based matching methods to incorporate spatial distance between units in addition to measured covariates, so that only units that are close in space can be matched to one another. Under causal assumptions like those in Section 3, this approach could mitigate unmeasured spatial confounding. Our paper can be seen as formalizing the assumptions under which including spatial coordinates in a propensity score or matching method is a valid identification strategy and as providing a nonparametric, doubly robust alternative for estimation.
Osama et al. (2019) consider a reparametrization of a spatially-varying coefficient model and propose a novel method for estimating the spatially-varying exposure coefficient without having to estimate $g(S)$. This methodology shares similarities with the spatial+ method of Dupont et al. (2020) but has the advantage of handling heterogeneous effects. However, the method deals only with linear exposure effects.
In the context of discrete, rather than continuous, spatial domains (i.e., areal data), Schnell and Papadogeorgou (2020) augment the traditional PLM with a parametric assumption about the joint distribution of the unmeasured spatial confounder and the exposure, which results in a likelihood model for unmeasured confounding. The paper also considers the estimation of non-linear effects by extending $X_{i}\beta$ to a general function $\eta(X_{i})$. However, the method relies on strong parametric identifying assumptions about the relationship between $U$ and $X$, and it is not clear when these would hold in practice or how practitioners would reason clearly about them, especially in settings where $U$ may be an unidentified latent confounder.
Finally, see Reich et al. (2021) for an overview of nascent methods for the analysis of effects under spatial interference, in which one unit’s exposure or outcome can affect others’ in a spatial pattern; we do not deal with the possibility of interference in the present work.
@@ 5.1 Doubly robust estimation
A doubly robust estimator is consistent for the estimand of interest if at least one of two nuisance models is correctly specified, typically a model for the propensity score and a model for the outcome regression. Doubly robust versions of the nonparametric outcome regression are available, e.g., Kennedy et al. (2017), but following the discussion in Section 3.4 we instead focus here on methods for doubly robust estimation of a shift intervention effect. Specifically, we describe the estimation of $\mu_{\delta}\equiv E[Y(X+\delta)]$, since the other term in the shift intervention contrast, $E[Y(X)]$, is trivially estimated by the sample average of $Y$. This estimand was first considered by Díaz Muñoz and van der Laan (2012) and Haneuse and Rotnitzky (2013), with additional theory established in Díaz et al. (2021). The following provides a presentation of the estimator from Díaz Muñoz and van der Laan (2012) for the purposes of exposition. In our simulations and data analysis we use a specific case of the estimator presented in Haneuse and Rotnitzky (2013) (see Supplement for details), which is asymptotically equivalent to the former, with the guarantee of the estimator lying in the parameter space (should the parameter space be bounded).
Let $\hat{m}$ be an estimator of $E[Y|X,S]$ and $\hat{\tau}$ be an estimator of $f(X|S)$. Consider the estimator
$\hat{\mu}_{\delta}=n^{-1}\sum_{i=1}^{n}\left[\frac{\hat{\tau}(X_{i}-\delta|S)}{\hat{\tau}(X_{i}|S_{i})}(Y_{i}-\hat{m}(X_{i},S_{i}))+\hat{m}(X_{i}+\delta,S_{i})\right]$
@@ 6 Simulations
We also simulate scenarios where no method is capable of returning consistent results due to scenarios causing violation of the identification assumptions such as noise in the confounding surface or smoothness of the exposure. In each case, the exposure is continuous; the causal contrast of interest is the effect of an intervention shifting the exposure by $+1$. More details of the simulation settings are provided in Supplementary Section S6.1, and code can be found at https://github.com/bjg345/causal_spatial.
@@
6.1 Spatially structured heterogeneous effect
@@
In the structural heterogeneous effect scenario (Figure 1(a)), only the flexible models and spline_interaction are correctly specified [^1] The exposure model of the DML methods misspecify the variances, but correctly specify the means.. The heterogeneity, since it is driven by the confounding variable, causes severe bias in all methods that assume a constant effect.
@@ 7 Application
For the present study, we use records of births from the California Department of Public Health’s Birth Statistical Master Files with mother’s home addresses geocoded using ArcGIS; we restrict our analysis to conceptions in the year 2013. We obtain daily PM2.5 levels at 1 kilometer resolution from the NASA Socioeconomic Data and Applications Center (Di et al., 2021) and average them over the course of pregnancy. To measure greenspace, we use peak normalized difference vegetation index (NDVI) measured at 250-m resolution and aggregated at the census block level from the year 2013 (Didan et al., 2015). In all our analyses, we include the potential confounding variables: race/ethnicity, parity, Kotelchuck index, maternal age, maternal education, and census tract-level per capita income from the 2010 American Community Survey (sourced from Manson et al. (2023)). We exclude individuals with missing data, as well as pregnancies with recorded gestational lengths less than 22 weeks or greater than 44 weeks, mothers with recorded ages less than 15 or greater than 49, and birthweights recorded below 500 grams. This leaves us with $n=457,439$ live births. Latitude and longitude were recorded in the Web Mercator projection, and all quantitative variables were normalized prior to model-fitting.
(a) Average daily PM2.5 concentrations by census tract in 2013 (data from Centers for Disease Control and Prevention – National Environmental Public Health Tracking Network and Census Bureau (2010)). Note that the exposure values used in our study were more granular, as described in the text.
(b) Plot of average daily PM2.5 concentration exposure in $\mu g/m^{3}$ and birthweight in grams.
Figure 2: PM2.5 exposure effect study.
@@
The exposure shift of interest is 0.1 $\mu g/m^{3}$. We choose this small shift due to a lack of positivity for larger shifts. Under linearity this effect can be multiplied by 10 in order to recover the implied average effect of a 0.1 $\mu g/m^{3}$ shift in exposure, and if the effect of exposure is approximately locally linear this rescaling is likely to provide a reasonable approximation. Because 0.1 $\mu g/m^{3}$ is too small of an effect to be of policy relevance we recommend rescaling. After rescaling the values we obtain may be compared to the results of linear models (by appropriate rescaling), but the effective interpretation of the coefficients of those linear models relies on the assumption of effect linearity which we do not make for non-linear models. Except for the linear models, 95% confidence intervals are by bootstrap standard errors. Our estimates are recorded in Table 1.
@@
Beginning with the linear effect models, we see that the effect of adjusting for NDVI is similar to the effect of adjusting for spatial location. This is consistent with the understanding that spatial location stands in for the unmeasured confounder NDVI. However, an assumption of linear effect may not be consistent with the data. As we can see from Figure 2, there appears to be a threshold effect where an effect only appears at high levels of exposure. The non-linear effect GAM models assume a homogeneous effect but do not assume the effect is linear. The estimates from these models are different from those of the linear effect models, but they display a similar pattern in which adjusting for spatial location and NDVI has a qualitatively similar effect. However, adjusting for spatial location decreases the magnitude of the estimated effect by more than adjusting for NDVI, suggesting that adjusting for space “picks up” additional confounding variables.

### v12

@@ 4.1 Partially linear models
Splines or other basis functions are used to model $g$. Alternatively, a linear mixed model (LMM)
@@
is used with the random effects $w(s_{i})$ typically modeled with Gaussian process (GP), independent of $X$, i.e., $w$ is a random function of space. The spatial effect is assumed to fully capture any spatial dependence, leaving the errors $\epsilon$ to be iid.
When used for spatial confounding, the (sometimes implicit) assumption is that the fixed or random function of space controls for unmeasured spatial confounding. We have identified three assumptions on the data-generating process that are commonly made in the existing spatial confounding literature:
  1. 1.
    There exists an unmeasured variable $U$ which is a fixed function of spatial location $S$, and $U$ influences both $X$ and $Y$. If $S$ is considered random, then $U$ is random as well, and so we may say that $U$ and $X$ are correlated. If $S$ is considered fixed, then $U$ is a nonrandom entity that is empirically associated with $X$.
  2. 2.
    There exists an unmeasured variable $U$ which is a random function of spatial location, and $U$ influences both $X$ and $Y$. In this case, $U$ is correlated with $X$ whether or not $S$ is random.
  3. 3.
    There exists an unmeasured variable $U$ which is a zero-mean random function of spatial location, and $U$ influences $Y$, but not $X$. In this case, $U$ is not a “confounder”; it merely introduces dependence among the outcomes, and naive estimators (like OLS) of the effect of $X$ on $Y$ are unbiased, although they may be inefficient or be associated with incorrect standard errors.
It is generally not obvious which scientific contexts justify the consideration of fixed or random effects for the locations in the data generation process. In what follows, we will mainly consider case 1; that is, we assume that there exists an unmeasured variable $U:=U(S)$, which is a fixed function of spatial location $S$ and influences both $X$ and $Y$. If the locations $S$ are considered to be sampled randomly, then $U=U(S)$ is random as well, and so we may say that $U$ and $X$ are correlated. If $S$ is considered fixed, then $U$ is a nonrandom entity that is empirically associated with $X$. Although in some cases the assumption of random sampling of locations may be unreasonable (e.g., if locations are taken on a regular lattice), the distinction between random and fixed sampling of locations may not be crucial since inference may proceed conditional on the selected locations, assuming they are ancillary to the parameters of interest. With minimal loss of generality, we will treat $S$ as random in what follows. Note that the assumption that $U$ is a fixed function of space in the true data generation process should not be confused with the modeling choice to regress $Y$ on a random function of space, e.g. using a Gaussian process regression, which may still be a valid estimation strategy (subject to assumptions) in the same way a fixed parameter can be modeled as a random quantity by assigning a prior in a Bayesian setting.
In spatial statistics literature, it is common to model spatial effects as random functions; we see a few reasons for this. First, physical variables are rarely simple or orderly functions of space, and the complexity of real-world confounding surfaces (that is, a confounding variable as a function of two-dimensional spatial location) may be reminiscent of subjective notions of randomness. And second, Gaussian processes, a particular class of random functions, are popular for their theoretical properties, and they are empirically successful in estimating fixed functions by formally modeling them as random variables. However, there is nothing inherently random about the true complex functions or surfaces in most applications. For example, using a Gaussian process to fit topographical elevation does not commit the researcher to any model of how that pattern of elevation came about, or to the view that topographical elevation is a random quantity. Deciding whether to model a given variable as stochastic may have more to do with generalization and targets of inference than scientific facts. If we wish to generalize to future iterations of a process that will involve a new confounding surface, then we must have a probabilistic model for this new generation.
We caution that conceiving confounding variables as random functions of space in the true data generation process may lead to spurious claims of estimator unbiasedness. That is, in reality, a confounding variable may exist which leads to estimation error (or conditional bias), but in imaginary re-generations of the confounding variable, the error would average out to zero (an absence of unconditional bias). We contend that the latter fact is mostly irrelevant for scientific purposes. We illustrate this with a toy example in the supplemental Section S4.
Conditional on the exposure $X$, the measured covariates $C$ (if any), and the locations $S$, we assume that the response $Y$ is independent and identically distributed ($iid$). Note that this assumption does not ignore spatial dependence; rather it assumes that any spatial dependence is captured via the function $U$ of the spatial locations of the observations. This is a typical assumption in many spatial regression models that include flexible functions of space in the regression function and assume that residuals are $iid$.
For simplicity, we ignore measured confounders $C$ in this section, but all of the methods discussed below can immediately accommodate measured covariates. In what follows in this section, we assume (3) as the data generation model while the analysis model can be (3) or (4) or other choices. Finally, note that some methods described below were developed in the context of areal data (spatial data where “location” refers to a region), though in this manuscript we are most interested in geostatistical data, in which locations are point-referenced.
@@
Ordinary least squares (OLS): The naive, unadjusted estimate of the effect of $X$ on $Y$ is given by the ordinary least squares (OLS) estimator, i.e., linear regression of $Y$ on $X$. This is the baseline to which spatial confounding proposals are often compared. It is unbiased assuming the random effect model (4) as the true data generation process, but see Section S4 of the supplement for a toy illustration of how claims of unbiasedness of the OLS estimator can be misleading. When data are generated from a process like (3) with a spatial effect, it is well-known that this procedure does not adjust for spatial confounding–under any of the operative definitions thereof–and is biased in most spatial settings. See Dupont et al. (2020) for a discussion and Gilbert et al. (2023) for a formal proof of the inconsistency of the OLS estimator in the context of spatial confounding.
@@
Gaussian Process regression: The LMM analysis model (4) with a GP prior for the random function $w$ calculates probabilities under the assumption that the random effect is independent of the exposure. At first glance, this would seem to rule out the possibility of addressing causal confounding. However, LMMs (4) with GP prior can consistently estimate $\beta$ when the data are generated from a PLM (Bickel and Kleijn, 2012; Yang et al., 2015). Therefore both spline models and GP regression are valid analysis models for estimating the effect of exposure when the data are generated from (3) and (5), which accommodates scenarios of spatial confounding.
Generalized least squares (GLS): GLS generalizes OLS to allow for correlated errors. GLS can also account for spatial confounding under (3) and (5), because it is equivalent to marginalizing out the LMM over the GP prior. The consistency of GLS under spatial confounding in the partial linear model is formally established in (Gilbert et al., 2023).
@@
Restricted spatial regression: Restricted spatial regression (RSR) (Hodges and Reich, 2010), considers an analysis model that is similar to the LMM in (4) but restricts $w$ to be orthogonal to the exposure $X$. Thus, including the spatial effects does not change the fixed effect estimate and $\bm{\hat{\beta}}_{RSR}=\bm{\hat{\beta}}_{OLS}$. The fact that the estimates from the mixed-effects model, like the GP regression model or the GLS estimator obtained from marginalizing out the LMM, are different from the OLS estimate was seen as a problem by Hodges and Reich (2010). This overlooks the fact that OLS is inconsistent for $\beta$ in the PLM under spatial confounding, as discussed above. The viewpoint of Hodges and Reich (2010) can perhaps be seen as stemming from assuming the LMM (4), having $w$ uncorrelated with $X$, as the data-generating process, in which case the OLS is consistent. In this scenario, the GLS estimate, accounting for spatial structure via the error term, will differ from the OLS estimate in finite samples, but both are unbiased estimators. Even assuming (4) as the true data-generating process, RSR has been criticized by Khan and Calder (2020) and Zimmerman and Hoef (2021) as producing anti-conservative intervals. However, while estimators (like GP regression and GLS) based on (4) as the analysis model can be consistent under spatial confounding, as discussed above, (4) itself should not be considered as the data generation process as it precludes any possibility of spatial confounding by modeling $w$ to be independent of $X$. We emphasize that with respect to the issue of causal confounding, RSR does not even attempt to supply an unbiased effect estimate; rather, it assumes there is no confounding bias. (See Gilbert et al., 2023 for further discussion of RSR, GLS, and GP regression.)
@@
Spatially varying coefficient model (SVC): Gelfand et al. (2003) maintains the assumption of a linear outcome regression as in (3) or (4) but allows the slope of the regression to vary by location, i.e., $\beta=\beta(S_{i})$. This SVC model thus relaxes the assumption of effect homogeneity and can account for confounding under spatially heterogeneous linear effects.
4.2 Limitations of linear models
@@
Under the assumption that data are generated from a partially linear model (3) (or an SVC) and the exposures are generated from (5), the spline- and GP-based methods above can be used to estimate causal effects under spatial confounding. To see why, we note that the assumption of (3) and (5) as the true data generation process implies the causal identification assumptions proposed in Section 3, and thus that controlling for a flexible function of space suffices to control for unmeasured spatial confounding. This is stated formally in the proposition below. The proof is provided in the Supplement.
Proposition 3.
@@
Recent work in spatial statistics extended the GP regression to allow for non-linear covariate effects fitted flexibly using random forests or boosting, thereby creating GLS versions of these machine learning procedures (Saha et al., 2021; Sigrist, 2020; Zhan and Datta, 2024). These extensions still use additive models, implying that the effect of exposure is homogeneous in space. In Section 5, we propose a class of methods that simultaneously relax both the linearity and homogeneity assumptions.
@@ 5.1 Doubly robust estimation
A doubly robust estimator is consistent for the estimand of interest if at least one of two nuisance models is correctly specified, typically a model for the propensity score and a model for the outcome regression. Doubly robust versions of the nonparametric outcome regression are available, e.g., Kennedy et al. (2017), but following the discussion in Section 3.4 we instead focus here on methods for doubly robust estimation of a shift intervention effect. Specifically, we describe the estimation of $\mu_{\delta}\equiv E[Y(X+\delta)]$, since the other term in the shift intervention contrast, $E[Y(X)]$, is trivially estimated by the sample average of $Y$. This estimand was first considered by Díaz Muñoz and van der Laan (2012) and Haneuse and Rotnitzky (2013), with additional theory established in Díaz et al. (2021). The following provides a presentation of the estimator from Díaz Muñoz and van der Laan (2012) for the purposes of exposition. In our simulations and data analysis, we use a specific case of the estimator presented in Haneuse and Rotnitzky (2013) (see Supplement for details), which is asymptotically equivalent to the former, with the guarantee of the estimator lying in the space of possible values of $\mu_{\delta}$ (should the parameter space be bounded).
Let $\hat{m}$ be an estimator of $E[Y|X,S]$ and $\hat{\tau}$ be an estimator of $f(X|S)$, conditional probability density of $X$ given $S$. Consider the estimator
$\hat{\mu}_{\delta}=n^{-1}\sum_{i=1}^{n}\bigg{[}\frac{\hat{\tau}(X_{i}-\delta|S)}{\hat{\tau}(X_{i}|S_{i})}(Y_{i}-\hat{m}(X_{i},S_{i}))\\ +\hat{m}(X_{i}+\delta,S_{i})\bigg{]}$
@@ 6 Simulations
We also simulate scenarios where no method is capable of returning consistent results due to scenarios causing a violation of the identification assumptions, such as noise in the confounding surface or smoothness of the exposure. In each case, the exposure is continuous; the causal contrast of interest is the effect of an intervention shifting the exposure by $+1$. More details of the simulation settings are provided in Supplementary Section S6.1 and code can be found at https://github.com/bjg345/causal_spatial .
@@
6.1 Linear spatially heterogeneous effect
@@
In this linear structural heterogeneous effect scenario (Figure 1(a)), only the flexible models and spline_interaction are correctly specified [^1] The exposure model of the DML methods misspecify the variances, but correctly specify the means.. The heterogeneity, since it is driven by the confounding variable, causes severe bias in all methods that assume a homogeneous effect.
@@ 7 Application
For the present study, we use records of births from the California Department of Public Health’s Birth Statistical Master Files with mother’s home addresses geocoded using ArcGIS; we restrict our analysis to conceptions in the year 2013. We obtain daily PM2.5 levels at 1-kilometer resolution from the NASA Socioeconomic Data and Applications Center (Di et al., 2021) and average them over the course of pregnancy. To measure greenspace, we use peak normalized difference vegetation index (NDVI) measured at 250-m resolution and aggregated at the census block level from the year 2013 (Didan et al., 2015). In all our analyses, we include the potential confounding variables: race/ethnicity, parity, Kotelchuck index, maternal age, maternal education, and census tract-level per capita income from the 2010 American Community Survey (sourced from Manson et al. (2023)). We exclude individuals with missing data, as well as pregnancies with recorded gestational lengths less than or equal to 22 weeks or greater than or equal to 44 weeks, mothers with recorded ages less than 15 or greater than 49, and birthweights recorded below 500 grams. This leaves us with $n\approx 450,000$ live births. Latitude and longitude were recorded in the Web Mercator projection, and all quantitative variables were normalized prior to model-fitting.
Figure 2: PM2.5 exposure concentrations by census tract in 2013 (data from Centers for Disease Control and Prevention – National Environmental Public Health Tracking Network and Census Bureau (2010)). Note that the exposure values used in our study were more granular, as described in the text.
@@
The exposure shift of interest is 0.1 $\mu g/m^{3}$. We choose this small shift due to a lack of positivity for larger shifts. Under linearity, this effect can be multiplied by 10 in order to recover the implied average effect of a 0.1 $\mu g/m^{3}$ shift in exposure, and if the effect of exposure is approximately locally linear, this rescaling is likely to provide a reasonable approximation. Because 0.1 $\mu g/m^{3}$ is too small of an effect to be of policy relevance we recommend rescaling. After rescaling the values we obtain may be compared to the results of linear models (by appropriate rescaling), but the effective interpretation of the coefficients of those linear models relies on the assumption of effect linearity which we do not make for non-linear models. Except for the linear models, 95% confidence intervals are by bootstrap standard errors. Our estimates are recorded in Table 1.
@@
Beginning with the linear effect models, we see that the effect of adjusting for NDVI is similar to the effect of adjusting for spatial location. This is consistent with the understanding that spatial location stands in for the unmeasured confounder NDVI. However, an assumption of linear effect may not be consistent with air pollution exposure (see, for example, Lin et al. (2022) and Miller et al. (2024)). As we can see from Figure 2, there appears to be a threshold effect where an effect only appears at high levels of exposure. The non-linear effect GAM models assume a homogeneous effect but do not assume the effect is linear. The estimates from these models are different from those of the linear effect models, but they display a similar pattern in which adjusting for spatial location and NDVI has a qualitatively similar effect. However, adjusting for spatial location decreases the magnitude of the estimated effect by more than adjusting for NDVI, suggesting that adjusting for space “picks up” additional confounding variables.